# Critical Points, Stability, and Basins of Attraction of Three Kuramoto Oscillators with Isosceles Triangle Network

### v1

@@ 2. $(2)$
Similarly, for $\Theta^{*}_{6}$, we have the following theorem.
Theorem 2.
Assume $K_{1}=-K_{2}<0$. If the initial configurations of system (1) satisfy
@@ 4 Conclusion
We have discussed theoretical findings on the stability and the attraction regions of critical points in Kuramoto oscillators, which relate to the nonlinear dynamics of three flames interacting in an isosceles configuration. Importantly, the parameter of coupling strength may assume negative values. The co-existence of stable oscillators was analyzed, providing adequate conditions for their attraction basins. These findings enhance our qualitative understanding of the complex behaviors in Kuramoto oscillators, even in a minimal system comprising only three nodes.

### v2

@@ 1. $(1)$
  2. $(2)$
    If the initial configurations of system (1) satisfy
@@ 2. $(2)$
Proof.
Part $(1)$: Based on Proposition 2 and equations $\theta_{1}=\tilde{\theta}_{1},{\theta}_{2}=\tilde{\theta}_{2}+\frac{2\pi}{3},{\theta}_{3}=\tilde{\theta}_{3}+\frac{\pi}{3},$ we obtain the first assertion. The proof of Part $(2)$ is similar to $(1)$.
@@ 4 Conclusion
We discussed theoretical findings on the stability and the attraction regions of critical points in Kuramoto oscillators, which relate to the nonlinear dynamics of three flames interacting in an isosceles configuration. The parameter of coupling strength can take negative values. The co-existence of stable oscillators was analyzed with the adequate conditions for their attraction basins. These findings enhance our qualitative understanding of the complex behaviors in Kuramoto oscillators, even in a minimal system comprising only three nodes.